OpenStudy (anonymous):

Check my Answer! Which equation would best help solve the following problem? The height of a triangle is 6 m more than its base. The area of the triangle is 56 m^2. Find the length of the base. A. \(56 = \frac{ 1 }{ 2 }(x)(x + 6)\) B. \(56 = \frac{ 1 }{ 2 }( x^2 + 6) \) C. \(56 = x(x + 6)\) D. \(56 = x^2 + 6\) My answer: A. or B.

OpenStudy (vera_ewing):

A = (1/2)bh = 56m^2 bh = 112 b(b+6m) = 112m^2 b^2 + 6b - 112 = 0 (b + 14)(b-8) = 0, b = 8m (tossing out negative solution)

OpenStudy (anonymous):

Sorry for taking so long, I had to type out the answer choices. @vera_ewing